OpenStudy (anonymous):

simplify: [sqrt{x-2}+(5/(sqrt{x-2}))]/(x-2)

mathslover (mathslover):

\[\LARGE{\sqrt{x-2}+\frac{5}{\sqrt{x-2}}/(x-2)}\] this is the question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

we can write (x-2) as root(x-2) *root(x-2) replace that and solve

mathslover (mathslover):

\[\LARGE{\sqrt{x-2}+\frac{5}{\sqrt{x-2}}/(x-2)}\] ok .. \[\LARGE{\sqrt{x-2}+\frac{5}{\sqrt{x-2}*(x-2)}}\]

mathslover (mathslover):

can u solve it now ..? @btnatyam

OpenStudy (anonymous):

do you have to make a common denominator with √x−2 ?

OpenStudy (anonymous):

@mathslover

mathslover (mathslover):

no @btnatyam make \[\sqrt{x-2}(x-2)\] as common denominator well have to go now ..sorry :(

OpenStudy (anonymous):

thanks!